# On-board Mission Replanning for Adaptive Cooperative Multi-Robot Systems

### v1

@@ III-A Analysis for small problem sizes
We also evaluate how performance scales with both problem complexity (Fig. 4b-c) and size (Fig. 4c). Of the two major modifications to mTSP (pink lines in Fig. 4b-c), the introduction of multiple start locations (orange lines) causes a larger negative impact on performance than incorporating variable task times (green lines). We suggest this is because the addition of new start location nodes causes a large increase in the number of possible solutions (as per equation 1), whereas varying task time only affects solution cost, not the number of solutions. In terms of scaling with problem size, we see an expected decrease in performance as the number of sub-tasks increases (Fig. 4c), but we note this trend is not strong - indeed it is only clearly visible if we analyse the proportion of highest-tier solutions $\hat{mt}_{Opt-P}<0.01$ as in (Fig. 4c).
III-B Generalization
@@
Figure 5: (a) Comparison of average mission times (mean across 30,000 test scenarios) achieved by GATR-General (purple lines) and GATR-Specific (dark blue) across varying numbers of tasks, agents and task discretization levels; both closely track NOpt-LKH3 performance (dashed grey lines) and are much smaller than randomly generated missions (represented by Med-S, solid grey lines). (b) Comparison of percentage increase in mission time of GATR-General (purple) and GATR-Specific (dark blue) relative to LKH3 as a function of increasing problem size (number of sub-tasks per agent). Coloured lines and shaded regions show mean and standard deviations respectively, with average across whole dataset indicated by coloured text. The normalized runtime $\hat{rt}$ of GATR-General is also shown against the right-hand y-axis by grey lines with circle markers (GPU runtimes) and star markers (Raspberry Pi runtimes).
@@
A key implementation challenge for neural solvers is their inability to generalize to different problem setups (see penultimate column, Table I). But generalisation is critical for replanning during a mission as the number of agents, tasks and other parameters will change throughout. This issue can be addressed by training a model on a diverse range of problem setups, but most existing implementations are constrained by fixed input graph sizes, including the RL4CO we adopt here, which supports varying numbers of agents but not tasks. To overcome this limitation, we calculate the maximum number of vertices required in the graph, i.e., $V_{max}=n_{max}\times\delta_{max}+m_{max}+1$. Based on the scenarios, we create graphs with $V=n\times\delta+m+1$ vertices and pad it with $V_{max}-V$ vertices. These extra nodes are masked during the cross-attention procedures and will not be selected as the next action as shown in Equation 3. The fixed-size matrices with custom paddings allow us to parallelize the data generation process while maintaining data diversity, hence enabling generalisation. To demonstrate model generalisation, we trained a single model (GATR-General) to solve problems with 1-6 agents, 1-6 tasks, and 1-4 task discretization levels. We tested GATR-General across 27 problem sizes (all combinations of 2/4/6 agents, 2/4/6 tasks and 1/2/4 discretization levels) and compared it against 27 specifically-trained models (GATR-Specific), one for each problem size. All models were also trained across three seeds. For fair comparision, we maintained the same network architecture to ensure similar level of RAM utilization. We set up test sets of 10,000 scenarios for each problem size with three seeds (30,000 total) as in [13], and evaluated GATR-General and GATR-Specific against NOpt-LKH3 and Med-S only; brute-force enumeration is not possible for the largest models considered here. We also compared runtime between GATR-General $rt_{GATR-General}$ and NOpt-LKH3 $rt_{NOpt-LKH3}$ on both a high-performance GPU (single-node 40GB A100 GPU) and on edge hardware (Raspberry Pi 5) using normalized runtime $\hat{rt}=\frac{rt_{GATR-General}}{rt_{NOpt-LKH3}}\times 100\%$.
@@
Overall, both GATR-General and GATR-Specific closely follow LKH3’s performance across a range of problem sizes (Fig. 5a). GATR-General and GATR-Specific achieve mission times that are only 7.1% and 2.6% larger than LKH3 respectively, and this gap does not show strong signs of widening with increasing workload per agent (Fig. 5b, purple and dark blue lines). Furthermore, GATR-General’s runtime is just 0.27–1.18% of LKH3’s runtime on edge hardware (e.g. 0.007s vs 2.7s at lower end); this fast speed is essential for real-time on-board replanning. Normalized runtime $\hat{rt}$ also decreases with increasing workload per agent, likely because GAT model runtime remains nearly constant regardless of problem size [15], whilst LKH3’s runtime scales with number of nodes. Execution on the Raspberry Pi is faster than on the GPU because there is limited room for parallelization when replans are only required infrequently and in isolation.
@@ IV Discussion
Our GATR framework enables effective and efficient replanning for cooperative multi-agent robotic systems, but it also has flexibility for future extension to incorporate additional key real-world capabilities; multi-objective planning, probabilistic consensus planning and anticipatory planning.
@@ IV-1 Managing competing aims - multi-objective planning
In this study, our GATR method is designed to minimise overall mission time only, which implies that there are no practical constraints on maximum overall mission time or on fuel usage for each robotic agent, and that all tasks are expected to be completed. However, these assumptions are not valid for many real-world scenarios, where the ability to consider and explore trade-offs between mission completion time, task completion rate (percentage of tasks completed), and agent attrition rate (percentage of agents lost, e.g. due to insufficient fuel for return to homebase) is important. Existing heuristic or meta-heuristic methods are not designed to handle such multi-objective planning, but our encoder-decoder GATR framework can easily support adaptation to incorporate these considerations in future work, primarily through modifying our model’s reward function to include terms for task completion rate and agent attrition rate, but also through modification of the model’s state space (e.g. adding elements to represent agent fuel levels, sensor functionality, actuator health), asymmetric edge weights and node costs (e.g. to represent wind or currents, which influence agent traversal and task completion times). These adaptations would also allow replanning for response to mission environment changes (e.g. reduction in mission time window due to incoming storm or imposition of a no-fly zone, change in travel times due to imcreased wind [6]) or more granular changes in system health (e.g. reduction of agent speed [33]), in addition to responding to the changes already demonstrated in this study; i.e. in tasking (changes in location, number, time-cost of tasks) and high-level system health (loss or addition of agents).
@@ IV-2 Managing current uncertainty - probabilistic consensus planning
To develop our GATR method in this study, we have assumed for simplicity that all robotic agents have the same perfect (i.e. zero uncertainty) knowledge of the environment state. Under these idealised conditions, all agents will generate the same replan for any given scenario; we therefore do not explicitly consider which agent generates any replan. But in real-world dynamic environments with limited communication, agents’ observations of the environment will be subject to uncertainty, and these observations and uncertainties will differ between agents [34]. Both of these challenges can be addressed through future incorporation of probabilistic Bayesian methods into GATR, which is straightforward for RL-based models [35] such as ours, but not for meta-heuristic or heuristic planning methods. The lightweight and fast nature of GATR makes it feasible to use such a probabilistic approach for consensus replanning, by generating replans from multiple robotic agents with differing beliefs. This consensus replanning could be achieved via a proposer-responder framework, in which any agent within a subset can temporarily assume the role of leader [36], propose a plan, and allow other agents to vote based on their own beliefs about the global state, merging multiple beliefs to reduce input uncertainties and achieving replanning in a semi-decentralized fashion. The combination of diverse beliefs and independent planning fosters ensemble learning, where collective intelligence can be used to enhance the overall planning quality [37]. Our GATR framework can therefore serve as a fundamental building block for development of a probabilistic-guided, reliable and resilient semi-decentralized replanner.
@@ IV-4 Conclusion
In this study we have defined the Cooperative Mission Replanning Problem (CMRP), which represents a key challenge for developing resilient, autonomous multi-robot systems, and have developed GATR, a new model to solve this problem effectively and efficiently. This work enables on-board replanning for autonomous multi-robot systems, improving their adaptability and resilience in dynamic, uncertain environments.
-A Graph Attention Model Overview
-A1 Attention Mechanism
Following [39], each attention layer consists of multi-head attention (MHA) and a feedforward layer. MHA attends to multiple subspaces, capturing relationships between elements. Queries, keys, and values are projected as $q^{\prime}=W_{q}q$, $k^{\prime}=W_{k}k$, and $v^{\prime}=W_{v}v$. We use $M=8$ attention heads with a hidden dimension $d_{h}=128$. The attention weights are computed as:
$A=\text{softmax}(\frac{qk}{\sqrt{d_{k}}}\cdot Mask)$ (3)
where $Mask$ prevents attention to certain positions. The output of each attention head is computed as $z_{i}=Av^{\prime}$, and the final outputs are concatenated and linearly projected.
-A2 Encoder
The encoder embeds input graphs for the RL agent. Vertex features comprise coordinates and time cost, represented as $\mathbf{x_{i}}=[x_{i},y_{i},t_{i}]$. These features are transformed into embeddings: $h_{i}^{(0)}=W\mathbf{x_{i}}+b$. These embeddings are processed through $N$ attention layers to generate latent vertex embeddings, and the graph embedding is computed as $\bar{h}^{(N)}=\frac{1}{n}\sum_{i=1}^{n}h_{i}^{(N)}$.
-A3 Decoder
The decoder selects sub-tasks sequentially. The environment state embedding is defined as $\mathbf{x_{i}}=[m_{remaining},t_{cur},t_{max},dist]$ and transformed into $h_{state_{t}}=Linear(\mathbf{x_{i}})$. The context embedding combines the graph embedding with the previous vertex embedding: $h_{(c)}^{(N)}=\bar{h}^{(N)}+Linear([h_{state},h_{\pi_{t-1}}^{(N)}])$. Masked attention prevents selecting visited sub-tasks or invalid moves. The probability of selecting a vertex is computed as: $p_{i}=\frac{e^{u_{i}}}{\sum_{i}e^{u_{j}}}$, where $u_{j}=-\infty$ if $j\in$ Mask; otherwise $u_{j}=C\cdot\tanh\left(\frac{q^{\prime}\cdot k^{\prime}_{j}}{\sqrt{d_{k}}}\right)$, with $C=10$.
-B Training
We train with a batch size of 5000 and a learning rate of $1\times 10^{-4}$ for up to 50 epochs with early stopping. Using 1,000,000 training and 100,000 validation samples, model training on an NVIDIA A100 GPU takes $\sim$1 hour.

### v2

@@ III-A Analysis for small problem sizes
where $mt_{benchmark}$ can refer to either mission time from the optimal benchmark ($mt_{Opt-P}$) or the near-optimal benchmark ($mt_{NOpt-LKH3}$) as desired. Here, $\hat{mt}_{benchmark}=0$ indicates our model produces a solution with the same mission time as the chosen benchmark (the optimal solution or a near-optimal solution from the LKH3 solver); and $\hat{mt}_{benchmark}=1$ indicates our model’s solution has the same mission time as an average randomly generated mission plan. As expected, optimality of the solution decreases as the problem complexity increases from mTSP (pink line in Fig. 4b) to the full CMRP (dark blue line). However, Fig. 4b shows that 89.3% of GATR solutions still achieve $\hat{mt}_{Opt-P}<0.1$ (normalized mission times within 10% of optimal), indicating consistently high performance despite the increased problem complexity over mTSP, and 12% solutions actually outperform our LKH3 heuristic benchmark (evidenced by the dashed line to the left of 0.00 in Fig. 4b). We also evaluate how performance scales with both problem complexity (Fig. 4b-c) and size (Fig. 4c). Of the two major modifications to mTSP (pink lines in Fig. 4b-c), the introduction of multiple start locations (orange lines) causes a larger negative impact on performance than incorporating variable task times (green lines). We suggest this is because the addition of new start location nodes causes a large increase in the number of possible solutions (as per equation 1), whereas varying task time only affects solution cost, not the number of solutions. In terms of scaling with problem size, we run additional experiments with varying number of sub-tasks, taking the test set from 300 to 4,500 runs per method. We see an expected decrease in performance as the number of sub-tasks increases (Fig. 4c), but we note this trend is not strong - indeed it is only clearly visible if we analyse the proportion of highest-tier solutions ($\hat{mt}_{Opt-P}<0.01$ in Fig. 4c).
III-B Generalization and scaling to larger problem sizes
@@
Figure 5: (a) Comparison of average mission times (mean across 30,000 test scenarios) achieved by GATR-General (purple lines) and GATR-Specific (dark blue) across varying numbers of tasks, agents and task discretization levels; both closely track NOpt-LKH3 performance (dashed grey lines) and are much smaller than randomly generated missions (represented by Med-S, solid grey lines). (b) Comparison of percentage increase in mission time of GATR-General (purple) and GATR-Specific (dark blue) relative to LKH3 as a function of increasing problem size (number of sub-tasks per agent). Colored lines and shaded regions show mean and standard deviations respectively, with average across whole dataset indicated by colored text.
@@
A key implementation challenge for neural solvers is their inability to generalize to different problem setups (see penultimate column, Table I). But generalization is critical for replanning during a mission as the number of agents, tasks and other parameters will change throughout. This issue can be addressed by training a model on a diverse range of problem setups, but most existing implementations are constrained by fixed input graph sizes, including RL4CO that we adopt here, which supports varying numbers of agents but not tasks. To overcome this limitation, we calculate the maximum number of vertices required in the graph, i.e., $V_{max}=n_{max}\times\delta_{max}+m_{max}+1$. Based on the scenario, we create a graph with $V=n\times\delta+m+1$ vertices and pad it with $V_{max}-V$ vertices. These extra nodes are masked during the cross-attention procedures and will not be selected as the next action. The fixed-size matrices with custom paddings allow us to parallelize the data generation process while maintaining data diversity, hence enabling generalization. To demonstrate model generalization, we trained a single model (GATR-General) to solve problems with 1-6 agents, 1-6 tasks, and 1-4 task discretization levels. We tested GATR-General across 27 problem sizes (all combinations of 2/4/6 agents, 2/4/6 tasks and 1/2/4 discretization levels) and compared it against 27 specifically-trained models (GATR-Specific), one for each problem size. All models were also trained across three seeds. For fair comparison, we maintained the same network architecture to ensure similar level of RAM utilization. We set up test sets of 10,000 scenarios for each problem size with three seeds (30,000 total) as in [13], and evaluated GATR-General and GATR-Specific against NOpt-LKH3 and Med-S only; brute-force enumeration is not possible for large problems. Even assuming our best-case runtime for evaluating a single solution ($3\times 10^{-5}\text{ s}$) from our simpler experiments (see previous Section), where it took $\sim$1 minute to obtain an optimal mission plan via enumeration, the largest problem size considered here (6 agents, 24 sub-tasks) would instead require $3\times 10^{-5}\text{ s}\times\frac{(24+6-1)!}{(6-1)!}\text{ possible solutions}=7\times 10^{16}\text{ years}$. We also compared inference runtimes between GATR-General  and NOpt-LKH3 on both a high-performance GPU (single-node 40GB A100 GPU) and on edge hardware (Raspberry Pi 5 8GB), both using the same configurations (software dependencies detailed in the Appendix). We trained all models on the GPU, then transferred the resulting checkpoint files to the Pi for inference.
TABLE II: Runtime of GATR-General and NOpt-LKH3
4 Agents 6 Agents
6 Sub-tasks 24 Sub-tasks 6 Sub-tasks 24 Sub-tasks
Machine Method Abs. Rel. Abs. Rel. Abs. Rel. Abs. Rel.
Pi GATR-General 0.0078 (1.18%) 0.0070 (0.33%) 0.0072 (0.67%) 0.0074 (0.27%)
NOpt-LKH3 0.7157 2.2074 1.1301 2.7467
GPU GATR-General 0.0624 (11.11%) 0.0633 (3.70%) 0.0726 (7.97%) 0.0709 (3.38%)
NOpt-LKH3 0.6138 1.7741 0.9642 2.159
Absolute (Abs.) times given in seconds, with relative (Rel.) times as % w.r.t. NOpt-LKH3.
Standard Errors typically $3\times 10^{-4}$ s for GATR-General and $3-7\times 10^{-3}$ s for NOpt-LKH3 .
@@ 6 Sub-tasks 24 Sub-tasks 6 Sub-tasks 24 Sub-tasks
Overall, both GATR-General and GATR-Specific closely follow LKH3’s performance across a range of problem sizes (Fig. 5a). GATR-General and GATR-Specific achieve mission times that are only 7.1% and 2.6% larger than LKH3 respectively, and this gap does not show strong signs of widening with increasing workload per agent (Fig. 5b, purple and dark blue lines). Furthermore, GATR-General’s runtime is just 0.27–1.18% of LKH3’s runtime on edge hardware (e.g. 0.007 s vs. 2.7 s at lower end, see first two rows in Table II); this fast speed is essential for real-time on-board replanning, especially when many individual replans may be required across multiple agents, e.g. for consensus-based planning (see Discussion). Table II (compare first two rows) shows GATR-General maintains consistent runtimes across varying numbers of sub-tasks from 6 to 24 (similar to [15]), whereas LKH3’s runtime scales poorly, increasing by roughly a factor of four. Moreover, execution on the Raspberry Pi is faster than on the GPU because there is limited room for parallelization when replans are only required in isolation.
@@ IV Discussion
Our GATR framework enables effective and efficient replanning for cooperative multi-agent robotic systems, but also has the major advantage of enabling future extension to incorporate additional key capabilities; multi-objective, probabilistic consensus, and anticipatory replanning.
@@ IV-1 Managing competing aims - multi-objective planning
In this study, our GATR method is designed to minimize overall mission time only, which implies that there are no practical constraints on maximum overall mission time or on fuel usage for each robotic agent, and that all tasks are expected to be completed. However, these assumptions are not valid for many real-world scenarios, where the ability to consider and explore trade-offs between mission completion time, task completion rate (percentage of tasks completed), and agent attrition rate (percentage of agents lost, e.g. due to insufficient fuel for return to homebase) is important. Existing heuristic or meta-heuristic methods are not designed to handle such multi-objective planning, but our encoder-decoder GATR framework can easily support adaptation to incorporate these considerations in future work, primarily through modifying our model’s reward function to include terms for task completion rate and agent attrition rate, but also through modification of the model’s state space (e.g. adding elements to represent agent fuel levels, sensor functionality, actuator health), asymmetric edge weights and node costs (e.g. to represent wind or currents, which influence agent traversal and task completion times). These adaptations would also allow replanning for response to mission environment changes (e.g. reduction in mission time window due to incoming storm or imposition of a no-fly zone, change in travel times due to increased wind [6]) or more granular changes in system health (e.g. reduction of agent speed [35]), in addition to responding to the changes already demonstrated in this study; i.e. in tasking (changes in location, number, time-cost of tasks) and high-level system health (loss or addition of agents).
@@ IV-2 Managing current uncertainty - probabilistic consensus planning
To develop our GATR method in this study, we have assumed for simplicity that all robotic agents have the same perfect (i.e. zero uncertainty) knowledge of the environment state. Under these idealized conditions, all agents will generate the same replan for any given scenario; we therefore do not explicitly consider which agent generates any replan. But in real-world dynamic environments with limited communication, agents’ observations of the environment will be subject to uncertainty, and these observations and uncertainties will differ between agents [36]. Both of these challenges can be addressed through future incorporation of probabilistic Bayesian methods into GATR, which is straightforward for RL-based models [37] such as ours, but not for meta-heuristic or heuristic planning methods. The lightweight and fast nature of GATR (runtime order of 0.07 s) makes it feasible to use such a probabilistic approach for consensus replanning, by generating replans from multiple robotic agents with differing beliefs. This consensus replanning could be achieved via a proposer-responder framework, in which any agent within a subset can temporarily assume the role of leader [38], propose a plan, and allow other agents to vote based on their own beliefs about the global state, merging multiple beliefs to reduce input uncertainties and achieving replanning in a semi-decentralized fashion. The combination of diverse beliefs and independent planning fosters ensemble learning, where collective intelligence can be used to enhance the overall planning quality [39]. Our GATR framework can therefore serve as a fundamental building block for development of a probabilistic-guided, reliable and resilient semi-decentralized replanner.
@@ IV-4 Conclusion
In this study we have defined the Cooperative Mission Replanning Problem (CMRP), which represents a key real-world challenge for developing resilient, autonomous multi-robot systems, and have developed GATR, a new model to solve this problem effectively and efficiently. This work enables on-board replanning for cooperative multi-robot systems for the first time, improving their adaptability and resilience in dynamic, uncertain environments.
Graph Attention Model Overview: Following [41], we use query/key dimensionality $d_{k}=\frac{\text{Hidden Dimensions}}{\text{Attention Heads}}=\frac{128}{8}=16$.
Encoder: The features of the input graph’s vertices (coordinates and time cost) are represented as $\mathbf{x_{i}}=[x_{i},y_{i},t_{i}]$. These are transformed into embeddings: $h_{i}^{(0)}=Linear(\mathbf{x_{i}})$. The embeddings $h_{i}^{(0)}$ are processed through $N$ attention layers to generate vertex embeddings $h_{i}^{(N)}$, and the graph embedding is computed as $\bar{h}^{(N)}=\frac{1}{n}\sum_{i=1}^{n}h_{i}^{(N)}$.
Decoder: The state embedding is defined as $\mathbf{s}=[m_{remaining},t_{cur},t_{max},dist]$ and transformed into $h_{state}=Linear(\mathbf{s})$. The context embedding combines the graph embedding with the previous vertex embedding: $h_{(c)}^{(N)}=[\bar{h}^{(N)},Linear([h_{state},h_{\text{pre. vertex}}^{(N)}])]$. The probability of selecting a vertex is computed as: $p_{i}=\frac{e^{u_{i}}}{\sum_{i}e^{u_{j}}}$, where $u_{j}=$-$\infty$ if $j\in$ Mask; otherwise $u_{j}=10\cdot\tanh\left(\frac{q^{\prime}\cdot k^{\prime}_{j}}{\sqrt{d_{k}}}\right)$, where queries $q^{\prime}$ and keys $k^{\prime}$ are products of learnable weights and embeddings.
Training: We train with a batch size of 5000 and a learning rate of $1\times 10^{-4}$ for up to 50 epochs with early stopping. Using 1,000,000 training and 100,000 validation samples, model training on an NVIDIA A100 GPU takes $\sim$1 hour. Major software dependencies include: Python 3.10.16, RL4CO 0.5.0.dev, TorchRL 0.6.0, PyTorch-Lightning 2.3.8.